OpenStudy (anonymous):

Zaineb is deciding whether to buy a stereo at $695 plus GST and PST (15%) now, or to invest the money to buy the stereo in a year. Her account pays 7.35% per annum, compounded monthly. If she can buy the stereo at the same price next year, how much would she save by investing the money?

OpenStudy (anonymous):

The answer is $60.76 but I don't know how..

OpenStudy (anonymous):

I tried P as the Principle invested: 695 x 0.30 = 208.50, i, as the interest rate per compounding period = 7.35% /100 = 0.0735/12 = 0.006125 and n, number of compounding periods = 1(12). P = 695(0.30)-208.50 i = 0.0735/12 = 0.006125 n = 12 A = P(1+i)^n = 695(1 + 0.006125)^12 = 224.35 Interest earned = Amount - Principle = 224.35 - 208.8 (is it?) = 15.55

OpenStudy (anonymous):

total cost of the stereo is (695)*.15 + 695 = 799.25

OpenStudy (anonymous):

I got it wrong D;

OpenStudy (anonymous):

I was going to work it through with you but it seems to me you are using the wrong number for the principle :D

OpenStudy (anonymous):

Ok so she is going to invest 799.25 cents thus we use the compound interest formula A = P(1 + (r/n))^(nt) A is what we are looking for P = 799.25 r = 7.35%/100 = 0.0735 n = 12 t = 12

OpenStudy (anonymous):

ohh.. D:

OpenStudy (anonymous):

how did you know when to use the compound interest formula? What was my formula called? o.o

OpenStudy (anonymous):

Maybe take a break :d seems like you have the math down just comprehension is a bit off :P

OpenStudy (anonymous):

your formula is the same as mine you are just using i to denote (r/n)

OpenStudy (anonymous):

ohhhh, i see.

OpenStudy (anonymous):

but it should be to the power of nt not just n

OpenStudy (anonymous):

yeah I have it on my notes too but my math textbook had me use that formula. Oh, i see.

OpenStudy (anonymous):

the time is also 12? not 1?

OpenStudy (anonymous):

because I multiplied 1 with 12, instead of 12 by 12.

OpenStudy (anonymous):

t = 1

OpenStudy (anonymous):

oh, and n = 12.

OpenStudy (anonymous):

the n is denoting that there are 12 compounds in a year and the t is stating that it is 1 year

OpenStudy (anonymous):

if you don't mind me asking what are compounds? .. :$ oh, i see.

OpenStudy (anonymous):

So you wouldn't really see the t as an exponent in this question because it is 1

OpenStudy (anonymous):

oh, okay.

OpenStudy (anonymous):

1 compound is when the interest is added to the total amount of money

OpenStudy (anonymous):

invested

OpenStudy (anonymous):

so the amount you invested is increased by 7.35% every compound

OpenStudy (anonymous):

ohh, i see.

OpenStudy (anonymous):

if the interest was 15% on 695 after 1 compound your annual or A would equal 799.25

OpenStudy (anonymous):

ohhh, okay.

OpenStudy (anonymous):

2 compounds it would be 799.25(.15) + 799.25 = A but yeah did you get the right answer?

OpenStudy (anonymous):

it's 919.14

OpenStudy (anonymous):

and I subtract that from?

OpenStudy (anonymous):

You would subtract it from the total cost of the stereo 799.25 doesn't seem right

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Ok so she is going to invest 799.25 cents thus we use the compound interest formula A = P(1 + (r/n))^(nt) A is what we are looking for P = 799.25 r = 7.35%/100 = 0.0735 n = 12 t = 1 799.25(1+(0.0735/12))^(12*1) =860.01 860.01-799.25 = 60.76

OpenStudy (anonymous):

ohh.. i forgot to do the power.

OpenStudy (anonymous):

Well good luck I hope I was a help

OpenStudy (anonymous):

thank you!!